I cannot find a formula to calculate this answer. I know how to get to the labor cost, but not the Overhead cost.

MC Qu. 182 A companys prime costs... A companys prime costs total $4,509,000 and its conversion costs total $5,510,000. If
Answer #1

Prime cost = Direct material+Direct labor

4509000 = 2011000+X

Direct labor = 2498000

Conversion Cost = Direct labor + Overhead

5510000 = 2498000+X

X(Overhead) = 5510000-2498000 = 3012000

So answer is b) $3012000